Banks charge customers fees of up to $35 per cheque for writing “bounced” cheques - that is, cheques that exceed the balance in the account. It has been estimated that processing bounced cheques costs a bank roughly $1.50 per cheque. Thus, the profit margin on a bounced cheque is very high. Realizing this, some banks process cheques from largest to smallest. By doing this, they maximize the number of cheques that bounce if a customer overdraws an account.
Antonio Freeman had a balance of $1,200 in his chequing account on a day when the bank received the following five cheques for processing against his account:
Cheque Number Amount:
3150, 3158, 3162, 3165 3169 $35, $1,510, $40, $550 $180

Assuming a $35 fee per cheque is assessed by the bank, how much fee revenue would the bank generate if it processed cheques (1) from largest to smallest, (2) from smallest to largest, and (3) in the order of the cheque numbers?
1
Expert's answer
2016-11-03T15:35:09-0400
1. $175 - the largest amount for the check exceeds the account balance, that is all bounced check (5 х $35)
2. $35 - one remains untreated check for $1510 (1200-35-40-180-550 = 395)
3. $140 - has four rough check $1510, $40.00, $550.00 and $180.00 (1200-35 =1165) (4 х $35)